Wheat is produced under perfectly competitive conditions. Individual wheat farmers have U-shaped, long-run average cost curves that reach a minimum average cost of $3 per bushel when 1,000 bushels are produced.

c) Given the new demand curve described in part (b), what will be the new long- run equilibrium? (That is, calculate market price, quantity of wheat produced, and the new equilibrium number of farms in this new situation
1
Expert's answer
2019-07-16T12:43:59-0400

The new long-run price is p=$ 3

Demand Q= 3,200,000- (200,000X3)= $2,600,000

Hence the number of firm J = 2,600,000/ 1000

J= 2,600 firms